Two dice are rolled. Let X be a random variable denoting the sum of the numbers


on the two dice.


i) Give the probability distribution of X


ii) Compute the expected value of X and its variance

1
Expert's answer
2022-04-27T10:58:16-0400

i)


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c}
 & 1 & 2 & 3 & 4 & 5 & 6 \\ \hline
 1 & 2 & 3 & 4 & 5 & 6 & 7 \\
 \hdashline
 2 & 3 & 4 & 5 & 6 & 7 & 8 \\
 \hdashline
 3 &4 & 5 & 6 & 7 & 8 & 9 \\
 \hdashline
 4 & 5 & 6 & 7 & 8 & 9 & 10 \\
 \hdashline
 5 & 6 & 7 & 8 & 9 & 10 & 11 \\
 \hdashline
 6 & 7 & 8 & 9 & 10 & 11 & 12 \\
 \hdashline
\end{array}"


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & p(x) \\ \hline
 2 & 1/36 \\
 \hdashline
 3 & 1/18 \\
 \hdashline
 4 & 1/12 \\
 \hdashline
 5 & 1/9 \\
 \hdashline
 6 & 5/36 \\
 \hdashline
 7 & 1/6 \\
 \hdashline
 8 & 5/36 \\
 \hdashline
 9 & 1/36 \\
 \hdashline
 10 & 1/12 \\
 \hdashline
 11 & 1/18 \\
 \hdashline
 12 & 1/36 \\
 \hdashline
\end{array}"

ii)


"E(X)=2(\dfrac{1}{36})+3(\dfrac{2}{36})+4(\dfrac{3}{36})+5(\dfrac{4}{36})"

"+6(\dfrac{5}{36})+7(\dfrac{6}{36})+8(\dfrac{5}{36})+9(\dfrac{4}{36})"

"+10(\dfrac{3}{36})+11(\dfrac{2}{36})+12(\dfrac{1}{36})=7"

"Var(X)=\sigma^2=E(X^2)-(E(X))^2"


"=2^2(\dfrac{1}{36})+3^2(\dfrac{2}{36})+4^2(\dfrac{3}{36})+5^2(\dfrac{4}{36})"

"+6^2(\dfrac{5}{36})+7^2(\dfrac{6}{36})+8^2(\dfrac{5}{36})+9^2(\dfrac{4}{36})"

"+10^2(\dfrac{3}{36})+11^2(\dfrac{2}{36})+12^2(\dfrac{1}{36})-7^2"

"=\dfrac{35}{6}"